A can do a piece of work in 24 days. If B is 60% more efficient than A, then find the number of days required by B to do the twice as large as the earlier work.

Grade:10

1 Answers

Sunil Raikwar
10 years ago
Let the whole work be one unit.
So work done by A in one day=1/24
Work done by B in one day=(1/24)+ 60% of (1/24)=(1/24)+(1/40)=8/120=1/15
So B can complete the whole work in 15 days.

Thanks & Regards
Sunil Raikwar
askIITians faculty